OpenStudy (anonymous):

ind the sum of the following infinite geometric series, if it exists. 2 + 6 + 18 + 54 +…

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

25,982 Does not exist 23,567 29,034

mathslover (mathslover):

What is the common ratio here ?

OpenStudy (anonymous):

dont know that

mathslover (mathslover):

Can you determine the pattern here, 2,6,18,54,... ?

OpenStudy (anonymous):

add 12

mathslover (mathslover):

Nope, as 18 + 20 is not equal to 54. See, they are being multiplied by 3.

OpenStudy (anonymous):

ok right!

mathslover (mathslover):

Right?

OpenStudy (zzr0ck3r):

3>1 thus dne

OpenStudy (noelgreco):

The sum would only exist if the value of r were less than 1. The r here is 3.

mathslover (mathslover):

Good. So, common ratio is 3. Now, formula for sum of infinite geometric series is a/(1-r) where r is common ratio and a is the first term First term is 2 common ratio is 3 so sum = 2/(1-3) = -1 (Thus does not exist as the sum is not negative here)

OpenStudy (anonymous):

we already know the ratio is 3

OpenStudy (zzr0ck3r):

a_i/a_j <1 for all I = j+1 this is necessary for convergence

OpenStudy (zzr0ck3r):

i = j + 1

mathslover (mathslover):

The tricks are already mentioned above by @zz0ck3r and @NoelGreco .

OpenStudy (anonymous):

ohhhh ok so the correct answer would be is does not exist

mathslover (mathslover):

Got it @chocodropa7 ?

mathslover (mathslover):

Yes.

OpenStudy (anonymous):

Got IT!

OpenStudy (zzr0ck3r):

correct

OpenStudy (anonymous):

thank you

mathslover (mathslover):

You're welcome choco. Best of Luck

OpenStudy (zzr0ck3r):

Oh sorry @mathslover my tablet is hard to read and I did not know you were walking him through it.

mathslover (mathslover):

:) No problem zzr0ck3r :) Its fine and even good to see.